# Force-current structure in Markovian open quantum systems and its applications: geometric housekeeping-excess decomposition and thermodynamic trade-off relations

### v1

@@ IV.1 Geometric housekeeping-excess decomposition
We first consider the geometric housekeeping-excess decomposition in the open quantum systems as an application of the developed framework. The housekeeping-excess decomposition, also known as the adiabatic-nonadiabatic decomposition, was first proposed to study irreversibility more precisely by dividing entropy production into distinct contributions [18]: The housekeeping part quantifies inevitable dissipation to keep the system out of equilibrium, while the excess one is the dissipation due to transient dynamics. It allows us to obtain tighter inequalities, focusing on physically distinct aspects of the dynamics [36, 37, 26, 27, 28, 29, 32, 30, 31]. In classical systems, essentially two concrete approaches have been proposed: one uses the long time limit of the dynamics (thus, the steady state) [19, 57], and the other focuses on the gradient structure of the dynamics [25, 32].
Some studies have revealed that we can establish such a decomposition for quantum systems [34, 35] based on the former strategy. However, their method requires a steady state that satisfies a particular condition so that the decomposed EPRs are positive. Although this problem also exists in the classical method [23, 24], it was recently overcome by using the geometric method, which involves the instantaneous gradient structure rather than the steady state [26, 27, 28, 29, 33, 30]. Still, the geometric decomposition has yet to be obtained for the quantum master equation; thus, a generally applicable decomposition of EPR is still lacking.
@@ IV.1.6 Unseparability
We discuss that the decomposed EPRs are not separable, unlike the total EPR; if we do the minimization (40) for each $k$, we will get a force of the form $\bigoplus_{k\in K}(-\nabla_{\mathbb{L}_{k}}\phi_{k})$, which is not a globally conservative force but a fully-locally conservative force. In fact, this fully local conservativeness is automatically satisfied if the dissipations are mediated by thermal baths; then, the expression in Eq. (15) implies $\mathbb{F}_{k}(\rho)=-\nabla_{\mathbb{L}_{k}}\phi_{k}(\rho)$, where $\phi_{k}(\rho)=\beta_{k}H+\ln\rho$. On the other hand, global conservativeness, namely, expression $\mathbb{F}(\rho)=-\nabla_{\mathbb{L}}\phi$ by a single potential operator $\phi$, is available only when the system is free from external driving so that it relaxes to the equilibrium. Therefore, in the decomposition, it is crucial that the minimization is implemented simultaneously to quantify the external effect on the dissipation.
Formally, we may define partial elements because the conservative force and current keep the direct-sum structure as $\mathbb{F}_{\mathrm{c}}(\rho)=\bigoplus_{k\in K}(-\nabla_{\mathbb{L}_{k}}\phi_{*}(\rho))$ and $\mathbb{J}_{\mathrm{c}}(\rho)=\bigoplus_{k\in K}\mathcal{M}_{\Gamma_{k}\otimes\rho}(-\nabla_{\mathbb{L}_{k}}\phi_{*}(\rho))$. However, in addition to the point that conservativeness is crucial only globally, we should be aware that the first equation in Eq. (42), $\nabla_{\mathbb{L}}\mathbb{J}_{\mathrm{c}}(\rho)=\mathcal{D}(\rho)$, does not mean that each current component $\mathcal{M}_{\Gamma_{k}\otimes\rho}(\nabla_{\mathbb{L}_{k}}\phi_{*}(\rho))$ gives $\mathcal{D}_{k}(\rho)$ when operated by $\nabla_{\mathbb{L}_{k}}$; therefore, the fragments of the housekeeping and excess EPRs would not be associated with conservativeness or minimum dissipation in the global sense. We note that, however, it could be interesting to consider partially-local conservativeness if we have a reasonable partition of $K$ into a few parts, as in the bipartite systems considered in information thermodynamics [64].
@@ IV.2.3 Time-dissipation uncertainty relation
If $X$ is time-independent and commutes with $H$ or $\rho$, then $\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho))=d\langle X\rangle/dt$ and the expectation value of $X$ approximately changes $\tau_{X}\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho))$ in the interval $\tau_{X}$. On the other hand, the “standard deviation” of the change would approximately be given by $\sqrt{2\mathscr{D}_{X}\tau_{X}}$ once we admit $\mathcal{D}_{X}$ is a quantum extension of diffusivity. By definition of $\tau_{X}$, we have
@@ V.1 Two-level system attached to two heat baths
Figure 3: Summary of the results in the first example. (a) We simulate the dissipative dynamics of a two-level system coupled to two heat baths at different temperatures. (b) We begin with a pure state $(|g\rangle+|e\rangle)/\sqrt{2}$ added a small noise to avoid divergence of $\ln\rho$. The coherence $\rho_{ge}$ vanished while the classical population converges to the stationary distribution. We do the simulation with parameters $\omega=1$, $\beta_{h}=0.58$, $\beta_{\ell}=1$, and $\gamma_{+}=\gamma_{-}=0.1$. (c) We plot the EPRs; as the system relaxes, the total and housekeeping EPRs converge to a single nonzero value, and the excess EPR becomes zero. The approximation formula reproduces the excess EPR well with a small error (shown in the inset).
@@
Now, the components of the force and current operators read
@@
where $\Delta$ is a real number corresponding to the potential difference between energy eigenstates, while $\psi$ are non-classical terms. As discussed in Sec. IV.1.5, such a $\phi$ is obtained by solving Eq. (54). For the present model, it is possible to solve the equation analytically. However, the exact form of the solution, given in Appendix E, is so complicated that it is little tractable. Instead, here we consider some extreme cases.
First, when the system has no coherence, $\rho_{ge}=\rho_{eg}=0$, we recover the classical result provided in Ref. [28] (sign is flipped since the “positive” direction of the jump is opposite),
@@ VI Conclusion
We would like to end this paper with some more comments on future directions: In this paper, we focused on the Euclidean geometry of forces. On the other hand, in classical systems, we can develop another type of theory that has more to do with optimization and the variational principle by using a non-Euclidean geometry called information geometry [29]. It would be an interesting avenue to consider the connection between thermodynamics and information geometry.
@@ Acknowledgements.
K.Y. thanks Ken Hiura and Naruo Ohga for their suggestive comments. K.Y. is supported by Grant-in-Aid for JSPS Fellows (Grant No. 22J21619). S.I. is supported by JSPS KAKENHI Grants No. 21H01560, No. 22H01141, No. 23H00467, and No. 24H00834, JST ERATO Grant No. JPMJER2302, and UTEC-UTokyo FSI Research Grant Program. This research was supported by JSR Fellowship, the University of Tokyo.
@@ Appendix A Mass action kinetics
$=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits_{\mathcal{H}}\big{(}(D^{*}(X^{2})-\{D^{*}(X),X\})\rho\big{)},$ (137)
@@
Next, we show the general inequality for Hermitian operators $B(>0)$ and $C$
@@
with $\phi_{a}\in\mathbb{R}$ and $\phi_{b}\in\mathbb{C}$. Here, note that the parametrization is different from the main text; since adding or subtracting a matrix proportional to the identity to $\phi$ does not affect the value of $\nabla_{\mathbb{L}}\phi$, $\phi_{a}$ equates to $-\Delta$ in Eq. (93). Further, $\phi_{b}$ corresponds to $\psi^{*}$. Consequently, we have
@@
Furthermore, we additionally assume $\rho$ is close to the steady state. $J_{\mathrm{cl}}(\rho)$ vanishes at the steady state, which can be confirmed by plugging Eq. (90) into the definition of $J_{\mathrm{cl}}(\rho)$. Thus, we can neglect terms related to $J_{\mathrm{cl}}(\rho)$ to get the approximation formulas
@@ Appendix F Numerical evidence of convergence
Figure 5: We compute the relative entropy between $\rho_{t}$ and the generalized Gibbs ensemble (GGE). The panels correspond to three values of the conserved quantity $Q$ (defined in Eq. (103)): $\langle Q\rangle=0.1,q^{\mathrm{eq}},$ and $0.7$, where $q^{\mathrm{eq}}=\mathop{\mathrm{tr}}\nolimits(Qe^{-\beta H}/Z)\simeq 0.197$. The GGE $\rho_{\beta,\lambda(\beta,q)}$ is chosen to have the same value of $q=\langle Q\rangle$ as the initial states. For each value of $q$, we prepare three initial states that have the designated value of $q$ and respectively close to $|gg\rangle$, $|+\rangle=(|ge\rangle+|eg\rangle)/\sqrt{2}$, or $|ee\rangle$. We observe exponential decay to the corresponding GGE in any case (though the convergence speed is diverse).

### v3

@@ IV.1 Geometric housekeeping-excess decomposition
We first consider the geometric housekeeping-excess decomposition in the open quantum systems as an application of the developed framework. The housekeeping-excess decomposition, also known as the adiabatic-nonadiabatic decomposition, was first proposed to study irreversibility more precisely by dividing entropy production into distinct contributions [21]: The housekeeping part quantifies inevitable dissipation to keep the system out of equilibrium, while the excess is dissipation due to transient dynamics. It allows us to obtain tighter inequalities, focusing on physically distinct aspects of the dynamics [40, 41, 29, 30, 31, 32, 33, 36, 34, 35, 33]. In classical systems, essentially two concrete approaches have been proposed: one uses the long time limit of the dynamics (thus, the steady state) [22, 65], and the other focuses on the gradient structure of the dynamics [28, 36].
Some studies have revealed that we can establish such a decomposition for quantum systems [38, 39] based on the former strategy. However, their method requires a steady state that satisfies a particular condition so that the decomposed EPRs are positive. Although this problem also exists in the classical method [26, 27], it was recently overcome by using the geometric method, which involves the instantaneous gradient structure rather than the steady state [29, 30, 31, 32, 33, 37, 34]. Still, the geometric decomposition has yet to be obtained for the quantum master equation; thus, a generally applicable decomposition of EPR is still lacking.
@@ IV.1.6 Unseparability
We discuss that the decomposed EPRs are not separable, unlike the total EPR; if we do the minimization (50) for each $k$, we will get a force of the form $\bigoplus_{k\in K}(-\nabla_{\mathbb{L}_{k}}\phi_{k})$, which is not a globally conservative force but a fully-locally conservative force. In fact, this fully local conservativeness is automatically satisfied if the dissipations are mediated by thermal baths; then, the expression in Eq. (15) implies $\mathbb{F}_{k}(\rho)=-\nabla_{\mathbb{L}_{k}}\phi_{k}(\rho)$, where $\phi_{k}(\rho)=\beta_{k}H+\ln\rho$. On the other hand, global conservativeness, namely, expression $\mathbb{F}(\rho)=-\nabla_{\mathbb{L}}\phi$ by a single potential operator $\phi$, is available only when the system is free from external driving so that it relaxes to the equilibrium. Therefore, in the decomposition, it is crucial that the minimization is implemented simultaneously to quantify the external effect on dissipation.
Formally, we may define partial elements because the conservative force and current keep the direct-sum structure as $\mathbb{F}_{\mathrm{c}}(\rho)=\bigoplus_{k\in K}(-\nabla_{\mathbb{L}_{k}}\phi_{*}(\rho))$ and $\mathbb{J}_{\mathrm{c}}(\rho)=\bigoplus_{k\in K}\mathcal{M}_{\Gamma_{k}\otimes\rho}(-\nabla_{\mathbb{L}_{k}}\phi_{*}(\rho))$. However, in addition to the point that conservativeness is crucial only globally, we should be aware that the first equation in Eq. (52), $\nabla_{\mathbb{L}}\mathbb{J}_{\mathrm{c}}(\rho)=\mathcal{D}(\rho)$, does not mean that each current component $\mathcal{M}_{\Gamma_{k}\otimes\rho}(\nabla_{\mathbb{L}_{k}}\phi_{*}(\rho))$ gives $\mathcal{D}_{k}(\rho)$ when operated by $\nabla_{\mathbb{L}_{k}}^{*}$; therefore, the fragments of the housekeeping and excess EPRs would not be associated with conservativeness or minimum dissipation in the global sense. We note that, however, it could be interesting to consider partially-local conservativeness if we have a reasonable partition of $K$ into a few parts, as in the bipartite systems considered in information thermodynamics [72].
@@ IV.2.3 Time-dissipation uncertainty relation
If $X$ is time-independent and commutes with $H$ or $\rho$, then $\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho))=d\langle X\rangle/dt$ and the expectation value of $X$ approximately changes $\tau_{X}\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho))$ in the interval $\tau_{X}$. On the other hand, the “standard deviation” of the change would approximately be given by $\sqrt{2\mathscr{D}_{X}\tau_{X}}$ once we admit $\mathscr{D}_{X}$ is a quantum extension of diffusivity. By definition of $\tau_{X}$, we have
@@ V.1 Two-level system attached to two heat baths
Figure 3: Summary of the results in the first example. (a) We simulate the dissipative dynamics of a two-level system coupled to two heat baths at different temperatures. (b) We begin with a pure state $(|g\rangle+|e\rangle)/\sqrt{2}$ added a small noise to avoid divergence of $\ln\rho$. The coherence $\rho_{ge}$ vanished while the classical population converges to the stationary distribution. We do the simulation with parameters $\omega=1$, $\beta_{h}=0.58$, $\beta_{\ell}=1$, and $\gamma_{+}=\gamma_{-}=0.1$. (c) We plot the EPRs. The excess and housekeeping EPRs are computed by solving Eq. (64).
As the system relaxes, the total and housekeeping EPRs converge to a single nonzero value, and the excess EPR becomes zero. The approximation formula reproduces the excess EPR well with a small error (shown in the inset).
@@
Now, the components of the force and current operators, defined in Eqs. (11) and (12), read
@@
where $\Delta$ is a real number corresponding to the potential difference between energy eigenstates, while $\psi$ are non-classical terms. As discussed in Sec. IV.1.5, such a $\phi$ is obtained by solving Eq. (64). The excess and housekeeping EPRs can be computed by using the potential as $\dot{\Sigma}^{\mathrm{hk}}=\lVert\mathbb{F}(\rho)-(-\nabla_{\mathbb{L}}\phi)\rVert_{\rho}^{2}$ and $\dot{\Sigma}^{\mathrm{ex}}=\lVert-\nabla_{\mathbb{L}}\phi\rVert_{\rho}^{2}$. For the present model, moreover, it is possible to solve the equation analytically. However, the exact form of the solution, given in Appendix E, is so complicated that it is little tractable. Instead, here we consider some extreme cases.
First, when the system has no coherence, $\rho_{ge}=\rho_{eg}=0$, we recover the classical result provided in Ref. [31] (the sign is flipped since the “positive” direction of the jump is opposite),
@@ VI Conclusion
We would like to end this paper with some more comments on future directions: In this paper, we focused on the Euclidean geometry of forces. On the other hand, in classical systems, we can develop another type of theory that has more to do with optimization and the variational principle by using a non-Euclidean geometry called information geometry [32, 33]. It would be an interesting avenue to consider the connection between thermodynamics and information geometry.
@@ Acknowledgements.
K.Y. thanks Ken Hiura, Naruo Ohga, and Artemy Kolchinsky for their suggestive comments. K.Y. is supported by Grant-in-Aid for JSPS Fellows (Grant No. 22J21619). S.I. is supported by JSPS KAKENHI Grants No. 21H01560, No. 22H01141, No. 23H00467, and No. 24H00834, JST ERATO Grant No. JPMJER2302, and UTEC-UTokyo FSI Research Grant Program. This research was supported by JSR Fellowship, the University of Tokyo.
@@ Appendix A Mass action kinetics
$=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits_{\mathcal{H}}\big{(}(\mathcal{D}^{*}(X^{2})-\{\mathcal{D}^{*}(X),X\})\rho\big{)},$ (147)
@@
Next, we show the general inequality for positive Hermitian operator $B$ and arbitrary operator $C$
@@
with $\phi_{a}\in\mathbb{R}$ and $\phi_{b}\in\mathbb{C}$. Here, note that the parametrization is different from the main text; $\phi_{a}$ can be identified with $-\Delta$ in Eq. (103) because adding or subtracting a matrix proportional to the identity to $\phi$ does not affect the value of $\nabla_{\mathbb{L}}\phi$. Further, $\phi_{b}$ corresponds to $\psi^{*}$. Consequently, we have
@@
Furthermore, we additionally assume $\rho$ is close to the steady state. $J_{\mathrm{cl}}(\rho)$ vanishes at the steady state, which can be confirmed by plugging Eq. (100) into the definition of $J_{\mathrm{cl}}(\rho)$. Thus, we can neglect second-order terms related to $J_{\mathrm{cl}}(\rho)$ to get the approximation formulas
@@ Appendix F Numerical evidence of convergence
Figure 5: We compute the relative entropy between $\rho_{t}$ and the generalized Gibbs ensemble (GGE). The panels correspond to three values of the conserved quantity $Q$ (defined in Eq. (113)): $\langle Q\rangle=0.1,q^{\mathrm{eq}},$ and $0.7$, where $q^{\mathrm{eq}}=\mathop{\mathrm{tr}}\nolimits(Qe^{-\beta H}/Z)\simeq 0.197$. The GGE $\rho_{\beta,\lambda(\beta,q)}$ is chosen to have the same value of $q=\langle Q\rangle$ as the initial states. For each value of $q$, we prepare three initial states that have the designated value of $q$ and respectively close to $|gg\rangle$, $|+\rangle=(|ge\rangle+|eg\rangle)/\sqrt{2}$, or $|ee\rangle$. We observe exponential decay to the corresponding GGE in any case (though the convergence speed is diverse). The irregular fluctuations at $10^{-15}$ are caused by floating-point arithmetic.